Formulate a system of equations for the situation below and solve. orchestra seats cost $70 apiece, and front balcony seats cost For the opening night at the Opera House, a total of 1000 tickets were sold. Front orchestra seats cost $90 aplece, rear orc $50 apiece. The combined number of tickets sold for the front orchestra and rear orchestra was 400 more than twice the num performance were $72,400. Determine how many tickets of each type were sold a total of 1000 tickets of front balcony tickets sold. The total receipts for the front orchestra rear orchestre front bakony
Answer #1

here let nmber of seats for front; rear orchestra and front balcony are x,y,z

hence x+y+z=1000

90x+70y+50z=72400

x+y-2z =400

solving above

x=320

y=480

z=200